IQ 8. Verify that the function =- √2² + y² + 2 is a solution of the three dimensional Laplace equation Uzz + Uyy + x = 0. 9. 8 and 9
Answer #1

8) u Jezy tzz Una otella (x2+y+z) 2(x+y++22354 Max = -51(+44 272_2.5(4+y+z= cho 1 [(x²+4² +2²) * = (x+y + z 293 [u²+ y + zAccording to the rules only one question will be answered